OpenStudy (anonymous):

Alex Stocks Up For Winter. He Buys 36 Cans Of Vegetables. He Pays 80 Cents Per Can For Tomatoes And 40 Cents Per Can For Corn, For A Total Cost Of 20.80 How Many Cans Of Tomatoes Does He Buy .

OpenStudy (anonymous):

@ChmE

OpenStudy (anonymous):

ok so we know there is a total of 36 cans

OpenStudy (anonymous):

This consists of t number of tomato and c number of corn cans

OpenStudy (anonymous):

Using this our first equation is t + c = 36

OpenStudy (anonymous):

Are you with me so far?

OpenStudy (anonymous):

Yeaa

OpenStudy (anonymous):

He pays 80 cents per tomato can so 0.80t, and 40 cents per corn so, 0.40c. And this equals a total of $20.80. So our second equation is 0.80t + 0.40c = 20.80

OpenStudy (anonymous):

Do you see that?

OpenStudy (anonymous):

now we have two equations for 2 unknowns and we can solve

OpenStudy (anonymous):

The question is only asking us for how many tomatoes, t.

OpenStudy (anonymous):

so we need to get one equation only in one variable, t. Your thoughts on how to do that??

OpenStudy (anonymous):

0.80t=20.80

OpenStudy (anonymous):

be careful... we can't just make c disappear. We need to enter a value in for c that will help us get the equation only in terms of t. Hint: solve the t + c = 36 for c. We can use this to help solve.

OpenStudy (anonymous):

What do you get if you solve the number of cans equation for c??

OpenStudy (anonymous):

t=36-c

OpenStudy (anonymous):

we need to solve it for c. which means c will be the variable by itself. Try again.

OpenStudy (anonymous):

we want c = something

OpenStudy (anonymous):

t=c

OpenStudy (anonymous):

t + c = 36 -t + t + c = 36 - t c = 36 - t

OpenStudy (anonymous):

do you see that?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Now we know what c is. This helps us because we can substitute it into the other equation to get it only in terms of t. This is called the substitution method. 0.80t + 0.40(36 - t) = 20.80

OpenStudy (anonymous):

Do you understand this step??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

distribute the 0.40 inside the () and tell me what you get

OpenStudy (anonymous):

So It Looks Like 0.80+(0.40+36-t)=20.80

OpenStudy (anonymous):

0.40(36 - t) is the same as 0.40 * (36 + -t). To do this we need to distribute the 0.40 to each term inside the ( ). This looks like this 0.40*36 + 0.40*-t

OpenStudy (anonymous):

Now type out the new equation

OpenStudy (anonymous):

Thanks For All The Help But I Have To Go Now ,

OpenStudy (anonymous):

ok. at this point we have only t so you can solve

OpenStudy (anonymous):

Good luck

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

I got 16 as my answer. now you can check it against ur own